# C-LoRA: Contextual Low-Rank Adaptation for Uncertainty Estimation in Large Language Models

### v2

@@ 3.4 Variational Bayesian Objective
$\nabla_{\varphi}\mathcal{L}(\mathbf{x},y)=\mathbb{E}_{\,\mathcal{E}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\nabla_{\boldsymbol{\varphi}}\left(\log p_{\boldsymbol{\theta}}(y|\mathbf{x},g_{\phi}(\mathcal{E},\mathbf{x}))-\log\frac{q_{\phi}(g_{\phi}(\mathcal{E},\mathbf{x})|\mathbf{x})}{p(g_{\phi}(\mathcal{E},\mathbf{x}))}\right)\right].$ (8)
3.5 Testing and Computational Complexity
@@
Test-time Sampling. Once C-LoRA training has converged, we obtain point estimate of the fine-tuned model by replacing $\mathbf{E_{x}}$ with its variational mean and computing $p(y|\mathbf{x},\boldsymbol{\mu}_{\mathbf{E}}(\mathbf{x}))$. As shown in Section 4, we denote these point estimate models by setting $M$ in our model names. Next, to demonstrate the quality of uncertainty estimation, we sample $\mathbf{E_{x}}$ from its inferred distribution $M$ times and, similar to (2), approximate the posterior predictive distribution using Monte Carlo sampling, yielding $p(y_{\mathrm{test}}|\mathbf{x}_{\mathrm{test}},\mathcal{D})=1/M\sum_{m=1}^{M}p(y|\mathbf{x},\mathbf{E}^{m}_{\mathbf{x}})$, where $\mathbf{E}^{m}_{\mathbf{x}}\sim q_{\phi}(\mathbf{E_{x}}|\mathbf{x})$. For example, for these calibrated models with $M=10$ drawn samples in Section 4, we label them with “M=10” in our model names. When “M=0”, the posterior mean is used directly for evaluation.
@@
Computational Complexity. In layer $l$, the contextual module is designed to model $q_{\phi}(\mathbf{E}^{l}_{\mathbf{x}}|\mathbf{x}^{l-1})$ where $\mathbf{x}^{l}$ is derived from natural language; hence, learning features from scratch becomes both non-trivial and computationally expensive. To mitigate this issue, we follow fan2021contextualdropoutefficientsampledependent and take advantage of the main model by feeding $\mathbf{z}^{l}$ into the contextual module instead. Additional computational cost of $\mathcal{O}(r^{4})$, stems from the auxiliary network $h^{l}_{\varphi}$, whose fully connected layers have size $C=\mathcal{O}(r^{2})\ll d$. This overhead is minimal compared to the main model’s per-layer operation cost of $\mathcal{O}(d^{2})$. As a result, we can efficiently estimate uncertainty at the sample level in a low-dimensional space, sidestepping the costly feature-learning burden within the contextual modules during fine-tuning.
@@ 4.1 Experimental Settings
Baselines. We provide performance comparisons of C-LoRA with well-established cutting-edge baselines that include Deep Ensemble balabanov2024uncertaintyquantificationfinetunedllms ; lakshminarayanan2017simplescalablepredictiveuncertainty ; wang2023loraensembleslargelanguage , Monte-Carlo dropout (MCD) pmlr-v48-gal16 , Bayesian LoRA with Backpropagation (BLoB) wang2024blobbayesianlowrankadaptation , a recent mean-field VI approach applied to LoRA parameters; Laplace-LoRA (LA) yang2024bayesianlowrankadaptationlarge , which uses a Laplace approximation over adapter weights; and Maximum A Posteriori (MAP), representing a deterministic point-estimate baseline.
LLM fine-tuning settings. Following prior work, we evaluate the performance of C-LoRA [^1] Using contextual modules comprising of two fully connected layers with 64 hidden units (C=64) each., by fine-tuning LLaMA2-7Btouvron2023llama2openfoundation using the PEFT librarypeft across six commonsense reasoning benchmarks, including both binary and multiple-choice classification tasks. For each task, the appropriate next-token logits are selected based on the format of the target output, which we include in the Appendix, and the model is trained to maximize the log-likelihood of the correct answer. Following previous works hu2021loralowrankadaptationlarge ; meo2024bayesianloralorabasedparameter ; wang2024blobbayesianlowrankadaptation , we apply LoRA to the query, value, and output projection layers using default hyperparameters. All models are fine-tuned with a batch size of 4. Baseline methods are trained for 5000 iterations, while C-LoRA is trained for 1500 iterations on all datasets, except for Winogrande-medium (WG-M), which is trained for 2000 iterations. To ensure consistent evaluation, we curate a training-validation split by reserving 80% of the original training set for fine-tuning and holding out the remaining 20% as a validation set. This validation split is used for early stopping and checkpointing based on a metric combining validation accuracy and calibration quality (see Appendix). Evaluation metrics. We report results using three key evaluation metrics: accuracy (ACC), expected calibration error (ECE), and negative log-likelihood (NLL). While ACC reflects raw predictive performance, ECE and NLL are widely adopted metrics for assessing the quality of uncertainty estimates. ECE measures the alignment between predicted confidence and actual correctness, while NLL evaluates the sharpness and correctness of the model’s predicted probability distribution. All reported results are averaged over three independent runs with different random seeds, and we report the mean $\pm$ standard deviation for each metric.
@@ 4.2 Evaluation Results under In-distribution Fine-tuning
In this section, we conduct our assessment of fine-tuning performances under the in-distribution scenario for six common-sense reasoning tasks. These tasks encompass: Winogrande-small (WG-S), Winogrande-medium (WG-M(sakaguchi2019winograndeadversarialwinogradschema, ), ARC-Challenge (ARC-C), ARC-Easy (ARC-E) clark2018thinksolvedquestionanswering , Open-BookQA (OBQA)mihaylov2018suitarmorconductelectricity , and BoolQ clark2019boolqexploringsurprisingdifficulty . For all the experiments, the same pre-trained LLaMA2-7B was used as the LLM backbone. For BLoB, we kept the same setting as in wang2024blobbayesianlowrankadaptation .
@@
Table 4.2 presents the performance of various uncertainty quantification methods applied to LoRA-tuned LLaMA2-7B across six common-sense reasoning datasets. Our proposed method, C-LoRA, as well as the baseline BloB, are evaluated under both deterministic (M=0) and stochastic (M=10) posterior sampling settings. While C-LoRA, does not achieve the highest accuracy, it maintains competitive performance across all datasets, with accuracies that are typically within 1–2% of the best methods (e.g., ARC-C: 67.79%, OBQA: 78.26%). Notably, C-LoRA (M=0)—which uses the posterior mean without sampling—performs similarly or slightly better than the sampled version on some tasks (e.g., ARC-C: 69.02% vs. 67.79%), suggesting that the contextual posterior mean alone offers a robust deterministic approximation.
@@
To further evaluate calibration, we apply post-hoc temperature scaling guo2017calibrationmodernneuralnetworks to both BLoB and C-LoRA and report the resulting ECE on six common-sense reasoning tasks in Table 4.2. Even before temperature scaling, C-LoRA (M=10) achieves stronger calibration than BLoB on nearly every dataset. After applying temperature scaling, C-LoRA (M=0, tmp) achieves the best or second-best ECE in 4 of the 6 datasets, including WG-S (4.00%), ARC-E (3.86%), ARC-C (6.30%), and WG-M (2.90%). Notably, it outperforms BLoB (M=10, tmp) in 5 out of the 6 datasets, despite using no posterior sampling. Interestingly, C-LoRA (M=10, tmp) does not always show improvement over its uncalibrated variant. On datasets like OBQA and BoolQ, temperature scaling can actually worsen ECE (e.g., from 4.00% to 6.27% on OBQA). This may be due to the fact that posterior sampling already flattens the predictive distribution, and further scaling can overcorrect, leading to underconfident predictions. These results suggest that C-LoRA (M=0) provides an effective balance between calibration quality and inference efficiency.
@@
MCD $2.50{\pm_{0.12}}$ $2.81\pm_{0.25}$ $1.13\pm_{0.04}$ $1.16\pm_{0.03}$ $1.01\pm_{0.07}$ $\underline{0.32}\pm_{0.00}$
Deep Ensemble $2.44{\pm_{0.23}}$ $2.20\pm_{0.03}$ $0.91\pm_{0.05}$ $1.04\pm_{0.09}$ $0.87\pm_{0.03}$ $\underline{0.32}\pm_{0.00}$
@@ 4.3 Robustness Under Distribution Shift
Table 3: Performance comparison on out-of-distribution datasets. The following results are evaluated using LLaMA2-7B fine-tuned on the OBQA dataset in Table 4.2.
@@
Table 4.3 presents model performance under both in-distribution (OBQA) and out-of-distribution (OOD) settings with smaller (ARC-E, ARC-C) and larger (Chem, Phy) distribution shifts. While C-LoRA does not achieve the highest accuracy under shift, it demonstrates strong robustness in uncertainty estimation, especially in calibration (ECE) and likelihood (NLL). In terms of accuracy, Deep Ensemble and BLoB generally maintain higher predictive performance across OOD datasets. However, C-LoRA (M=10) remains competitive, achieving 65.09% accuracy on ARC-C and 41.00% on Chem—nearly matching the highest-performing methods. While C-LoRA (M=10) sees a slight drop in accuracy under severe shift (e.g., Phy: 31.00%), it significantly outperforms all baselines in ECE and NLL, indicating more reliable and calibrated uncertainty estimates.
@@ 4.4 Ablation Study: Impact of Contextual Module
We also conduct the ablation study investigating the importance of the contextual module for overall performance. Since in C-LoRA we focus on contextualizing the distribution of matrix $\mathbf{E}\in\mathbb{R}^{r\times r}$, to examine the effect of the contextual module, we compare with the results of fine-tuning the model on different tasks using lightweight factorization introduced in Section 3.1, with mean-field variational inference without the contextual module; we refer to this setting as FE in Tables 4.4 and 4.4.
Table 4.4 underscores the importance of the contextual module by comparing FE and C-LoRA across the six common-sense reasoning tasks under in-distribution scenario. C-LoRA offers significant improvements with respect to ECE across almost all tasks for inference-time sample with both M=0 and M=10, except for BoolQ where it achieves the second best performance. Also, C-LoRA gives the lowest NLL across all tasks.
Furthermore, we assess the effect of the contextual module on generalization. Table 4.4 provides the performance comparison of the corresponding models fine-tuned on OBQA in Table 4.4 under similar distribution shifts as in Table 4.3. C-LoRA outperforms FE across all datasets in both smaller and larger distribution shifts, with respect to ECE. In addition, C-LoRA offers the lowest NLL on all tasks in both distribution shifts. With respect to accuracy, in all tasks C-LoRA achieves a comparable performance, except for Chem where it has a subpar performance; however, this shortcoming is a result of the tradeoff between a generalized model with a proper uncertainty quantification, and an overconfident model with a poor uncertainty estimation. In summary, these results indicate the significance of our auxiliary contextual module, particularly on enhancing the uncertainty quantification and generalization while maintaining competitive predictive performance.
@@
[…]
@@
Table 5: Impact of the contextual module on generalization. The performance comparison is conducted on out-of-distribution datasets using LLaMA2-7B fine-tuned on OBQA dataset in Table 4.4.

### v3

@@ 3.3 Amortized Inference for Contextual LoRA
$\mathcal{L}=\sum_{i=1}^{N}\left[\mathbb{E}_{\,q_{\boldsymbol{\phi}}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i})}\log p_{\boldsymbol{\theta}}(y_{i}|\mathbf{x}_{i},\mathbf{E}_{\mathbf{x}_{i}})-\mathrm{KL}(q_{\boldsymbol{\phi}}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i})\,\|\,p(\mathbf{E}_{\mathbf{x}_{i}}))\right].$ (6)
Here, $\boldsymbol{\theta}$ denotes the parameters of all $\mathbf{B}$ and $\mathbf{A}$ adapters across layers. Hence, $\boldsymbol{\phi}=\{\boldsymbol{\theta},\boldsymbol{\varphi}\}$ denotes all model parameters. Unlike standard variational inference as in (3), our formulation makes the distribution of $\mathbf{E}_{\mathbf{x}_{i}}$ input-dependent and replaces the single KL term with a sum of $N$ per-sample $\mathrm{KL}$ divergences, whose combined impact scales with $N$. With the objective defined in (6), we adopt a simple and fixed Gaussian prior $p(\mathbf{E_{x}})$ for learning $\mathbf{E_{x}}$ in each layer. The complete learning objective can be equivalently expressed as a summation over all samples: $\mathcal{L}=\sum_{(\mathbf{x},y)\in\mathcal{D}}\mathcal{L}(\mathbf{x},y)$. Then, we learn the deterministic parameters $\mathbf{B}$ and $\mathbf{A}$ together represented by $\boldsymbol{\theta}$ using the expected negative log-likelihood (NLL) term while excluding the $\mathrm{KL}$ term. This ensures that learning of $\boldsymbol{\theta}$ remains supervised, yielding less noisy gradients per sample computed as
$\nabla_{\boldsymbol{\theta}}\mathcal{L}(\mathbf{x},y)=\mathbb{E}_{\,q_{\boldsymbol{\phi}}(\mathbf{E}_{\mathbf{x}}|\mathbf{x})}\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(y|\mathbf{x},\mathbf{E}_{\mathbf{x}}).$ (7)
We approximate this expectation with a single Monte Carlo draw $\mathbf{E_{x}}\sim q_{\boldsymbol{\phi}}(\mathbf{E_{x}}|\mathbf{x})$ for each $\mathbf{x}$ sample.
Next, to update the auxiliary network parameters $\boldsymbol{\varphi}$, which appear in both the NLL and $\mathrm{KL}$ terms of (6), we apply the reparameterization trick kingma2015repara_trick to handle random sampling of $\mathbf{E_{x}}$. Concretely, in each layer, $\mathbf{E}^{l}_{\mathbf{x}}\sim\mathcal{N}(\boldsymbol{\mu}^{l}_{\mathbf{E}},{\boldsymbol{\Omega}^{l}}^{2}_{\mathbf{E}})$ is reparameterized as $\mathbf{E}^{l}_{\mathbf{x}}=\boldsymbol{\mu}^{l}_{\mathbf{E}}+\boldsymbol{\Omega}^{l}_{\mathbf{E}}\odot\mathcal{E}^{l}$, where $\mathcal{E}^{l}\in\mathbb{R}^{r\times r}$ is sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. Thus, sampling $\mathbf{E}^{l}_{\mathbf{x}}$ from $q_{\phi}(\mathbf{E}_{\mathbf{x}}|\mathbf{x})$ is equivalent to evaluating a deterministic mapping $g_{\phi}(\mathcal{E},\mathbf{x})$, enabling gradient computation via
$\nabla_{\boldsymbol{\varphi}}\mathcal{L}(\mathbf{x},y)=\mathbb{E}_{\,\mathcal{E}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\nabla_{\boldsymbol{\varphi}}\left(\log p_{\boldsymbol{\theta}}(y|\mathbf{x},g_{\boldsymbol{\phi}}(\mathcal{E},\mathbf{x}))-\log\frac{q_{\boldsymbol{\phi}}(g_{\boldsymbol{\phi}}(\mathcal{E},\mathbf{x})|\mathbf{x})}{p(g_{\boldsymbol{\phi}}(\mathcal{E},\mathbf{x}))}\right)\right].$ (8)
3.4 Posterior Predictive Inference and Model Complexity
@@
Posterior predictive inference and uncertainty estimation. Once C-LoRA training has converged, we obtain point estimate of the fine-tuned model by replacing $\mathbf{E_{x}}$ with its variational mean and computing $p(y|\mathbf{x},\boldsymbol{\mu}_{\mathbf{E}}(\mathbf{x}))$. As shown in Section 4, we denote these point estimate models by setting $M$ in our model names. Next, to demonstrate the quality of uncertainty estimation, we sample $\mathbf{E_{x}}$ from its inferred distribution $M$ times and, similar to (2), approximate the posterior predictive distribution using Monte Carlo sampling, yielding $p(y_{\mathrm{test}}|\mathbf{x}_{\mathrm{test}},\mathcal{D})=1/M\sum_{m=1}^{M}p(y|\mathbf{x},\mathbf{E}^{m}_{\mathbf{x}})$, where $\mathbf{E}^{m}_{\mathbf{x}}\sim q_{\boldsymbol{\phi}}(\mathbf{E_{x}}|\mathbf{x})$. For example, for these calibrated models with $M=10$ drawn samples in Section 4, we label them with “M=10” in our model names. When “M=0”, the posterior mean is used directly for evaluation.
@@
Reducing contextual module complexity through feature reuse. In layer $l$, the contextual module is designed to model $q_{\boldsymbol{\phi}}(\mathbf{E}^{l}_{\mathbf{x}}|\mathbf{x}^{l-1})$ where $\mathbf{x}^{l}$ is derived from natural language; hence, learning features from scratch becomes both non-trivial and computationally expensive. To mitigate this issue, we follow fan2021contextual and take advantage of the main model by feeding $\mathbf{z}^{l}$ into the contextual module instead. Additional computational cost of $\mathcal{O}(r^{4})$, stems from the auxiliary network $h^{l}_{\varphi}$, whose fully connected layers have size $C=\mathcal{O}(r^{2})\ll d$. This overhead is minimal compared to the main model’s per-layer operation cost of $\mathcal{O}(d^{2})$. As a result, we can efficiently estimate uncertainty at the sample level in a low-dimensional space, sidestepping the costly feature-learning burden within the contextual modules during fine-tuning.
@@ 4.1 Experimental Settings
Baselines. We provide performance comparisons of C-LoRA with well-established cutting-edge baselines that include Deep Ensemble balabanov2025uncertainty ; NIPS2017_9ef2ed4b ; wang2023loraensembleslargelanguage , Monte-Carlo dropout (MCD) pmlr-v48-gal16 , Bayesian LoRA with Backpropagation (BLoB) wang2024blobbayesianlowrankadaptation , a recent mean-field VI approach applied to LoRA parameters; Laplace-LoRA (LA) yang2024bayesianlowrankadaptationlarge , which uses a Laplace approximation over adapter weights; and Maximum A Posteriori (MAP), representing a deterministic point-estimate baseline.
LLM fine-tuning settings. Following prior work, we evaluate the performance of C-LoRA [^1] Using contextual modules comprising of two fully connected layers with 64 hidden units (C=64) each., by fine-tuning LLaMA2-7Btouvron2023llama2openfoundation using the PEFT librarypeft across six commonsense reasoning benchmarks, including both binary and multiple-choice classification tasks. For each task, the appropriate next-token logits are selected based on the format of the target output, which we include in the Appendix A, and the model is trained to maximize the log-likelihood of the correct answer. Following previous works hu2022lora ; meo2024bayesianlora ; wang2024blobbayesianlowrankadaptation , we apply LoRA to the query, value, and output projection layers using default hyperparameters. All models are fine-tuned with a batch size of 4. Baseline methods are trained for 5000 iterations, while C-LoRA is trained for 1500 iterations on all datasets, except for Winogrande-medium (WG-M), which is trained for 2000 iterations. To ensure consistent evaluation, we curate a training-validation split by reserving 80% of the original training set for fine-tuning and holding out the remaining 20% as a validation set. This validation split is used for early stopping and checkpointing based on a metric combining validation accuracy and calibration quality (see Appendix F).
Evaluation metrics. We report results using three key evaluation metrics: accuracy (ACC), expected calibration error (ECE), and negative log-likelihood (NLL). While ACC reflects raw predictive performance, ECE and NLL are widely adopted metrics for assessing the quality of uncertainty estimates. ECE measures the alignment between predicted confidence and actual correctness, while NLL evaluates the sharpness and correctness of the model’s predicted probability distribution. All reported results are averaged over three independent runs with different random seeds, and we report the mean $\pm$ standard deviation for each metric.
@@ 4.2 Evaluation Results under In-distribution Fine-tuning
In this section, we conduct our assessment of fine-tuning performances under the in-distribution scenario for six common-sense reasoning tasks. These tasks encompass: Winogrande-small (WG-S), Winogrande-medium (WG-M(10.1145/3474381, ), ARC-Challenge (ARC-C), ARC-Easy (ARC-E) clark2018thinksolvedquestionanswering , Open-BookQA (OBQA)mihaylov-etal-2018-suit , and BoolQ clark-etal-2019-boolq . For all the experiments, the same pre-trained LLaMA2-7B was used as the LLM backbone. For BLoB, we kept the same setting as in wang2024blobbayesianlowrankadaptation .
@@
Table 1 presents the performance of various uncertainty quantification methods applied to LoRA-tuned LLaMA2-7B across six common-sense reasoning datasets. As a complement, Figure 2 in Appendix K visualizes the results with bar plots. In this set of experiments, our proposed method, C-LoRA, as well as the baseline BLoB, are evaluated under both deterministic (M=0) and stochastic (M=10) posterior sampling settings. While C-LoRA, does not achieve the highest accuracy, it maintains competitive performance across all datasets, with accuracies that are typically within 1–2% of the best methods (e.g., ARC-C: 67.79%, OBQA: 78.26%). Notably, C-LoRA (M=0)—which uses the posterior mean without sampling—performs similarly or slightly better than the sampled version on some tasks (e.g., ARC-C: 69.02% vs. 67.79%), suggesting that the contextual posterior mean alone offers a robust deterministic approximation.
@@
To further evaluate calibration, we apply post-hoc temperature scaling pmlr-v70-guo17a to both BLoB and C-LoRA and report the resulting ECE on six common-sense reasoning tasks in Table 2. Bar plots of these results are presented in Figure 3 in Appendix K. Even before temperature scaling, C-LoRA (M=10) achieves stronger calibration than BLoB on nearly every dataset. After applying temperature scaling, C-LoRA (M=0, tmp) achieves the best or second-best ECE in 4 of the 6 datasets, including WG-S (4.00%), ARC-E (3.86%), ARC-C (6.30%), and WG-M (2.90%). Notably, it outperforms BLoB (M=10, tmp) in 5 out of the 6 datasets, despite using no posterior sampling. Interestingly, C-LoRA (M=10, tmp) does not always show improvement over its uncalibrated variant. On datasets like OBQA and BoolQ, temperature scaling can actually worsen ECE (e.g., from 4.00% to 6.27% on OBQA). This may be due to the fact that posterior sampling already flattens the predictive distribution, and further scaling can overcorrect, leading to underconfident predictions. Also, temperature scaling minimizes NLL, a proper scoring rule, but it does not directly optimize ECE, which is a binned, non-differentiable, and improper metric. Consequently, if a model is already reasonably well-calibrated or if its logit distribution exhibits certain local structures, temperature scaling may shift predictions in ways that degrade ECE in specific bins. Similar observations have been reported in prior work NEURIPS2019_8ca01ea9 , where temperature scaling was found to degrade classwise ECE. These results suggest that C-LoRA (M=0) provides an effective balance between calibration quality and inference efficiency.
@@
MCD $2.50{\pm_{0.12}}$ $2.81\pm_{0.25}$ $13\pm_{0.04} 1.1$$16\pm_{0.03}$ $1.01\pm_{0.07}$ $\underline{0.32}\pm_{0.00}$
Deep Ensemble $44{\pm_{0.23}} 2.2$$20\pm_{0.03}$ $0.91\pm_{0.05}$ $1.04\pm_{0.09}$ $0.87\pm_{0.03}$ $\underline{0.32}\pm_{0.00}$
@@ 4.3 Robustness Under Distribution Shift
Table 3: Performance comparison on out-of-distribution datasets. The following results are evaluated using LLaMA2-7B fine-tuned on the OBQA dataset in Table 1.
@@
Table 3 presents model performance under both in-distribution (OBQA) and out-of-distribution (OOD) settings with smaller (ARC-E, ARC-C) and larger (Chem, Phy) distribution shifts. Figure 4 in Appendix K illustrates bar plots of the results. While C-LoRA does not achieve the highest accuracy under shift, it demonstrates strong robustness in uncertainty estimation, especially in calibration (ECE) and likelihood (NLL). In terms of accuracy, Deep Ensemble and BLoB generally maintain higher predictive performance across OOD datasets. However, C-LoRA (M=10) remains competitive, achieving 65.09% accuracy on ARC-C and 41.00% on Chem—nearly matching the highest-performing methods. While C-LoRA (M=10) sees a slight drop in accuracy under severe shift (e.g., Phy: 31.00%), it significantly outperforms all baselines in ECE and NLL, indicating more reliable and calibrated uncertainty estimates.
@@ 4.4 Ablation Study: Impact of Contextual Module
We also conduct the ablation study investigating the importance of the contextual module for overall performance. Since in C-LoRA we focus on contextualizing the distribution of matrix $\mathbf{E}\in\mathbb{R}^{r\times r}$, to examine the effect of the contextual module, we compare with the results of fine-tuning the model on different tasks using lightweight factorization introduced in Section 3.1, with mean-field variational inference without the contextual module; we refer to this setting as FE in Tables 4 and 5.
Table 4 underscores the importance of the contextual module by comparing FE and C-LoRA across the six common-sense reasoning tasks under in-distribution scenario. C-LoRA offers significant improvements with respect to ECE across almost all tasks for inference-time sample with both M=0 and M=10, except for BoolQ where it achieves the second best performance. Also, C-LoRA gives the lowest NLL across all tasks.
Furthermore, we assess the effect of the contextual module on generalization. Table 5 provides the performance comparison of the corresponding models fine-tuned on OBQA in Table 4 under similar distribution shifts as in Table 3. C-LoRA outperforms FE across all datasets in both smaller and larger distribution shifts, with respect to ECE. In addition, C-LoRA offers the lowest NLL on all tasks in both distribution shifts. With respect to accuracy, in all tasks C-LoRA achieves a comparable performance, except for Chem where it has a subpar performance; however, this shortcoming is a result of the tradeoff between a generalized model with a proper uncertainty quantification, and an overconfident model with a poor uncertainty estimation. In summary, these results indicate the significance of our auxiliary contextual module, particularly on enhancing the uncertainty quantification and generalization while maintaining competitive predictive performance.
@@
C-LoRA (M=0) $\underline{7.16}{\pm_{2.92}}$ $\underline{12.28}\pm_{1.01}$ $\underline{5.75}\pm_{1.00}$ $\underline{6.07}\pm_{4.85}$ $\underline{5.10}\pm_{1.36}$ $\underline{1.46}\pm_{0.85}$
@@
Table 5: Impact of the contextual module on generalization. The performance comparison is conducted on out-of-distribution datasets using LLaMA2-7B fine-tuned on OBQA dataset in Table 4.